# MMM: Multilingual Mutual Reinforcement Effect Mix Datasets & Test with Open-domain Information Extraction Large Language Models

### v1

@@ Abstract
The Mutual Reinforcement Effect (MRE) represents a promising avenue in information extraction and multitasking research. Nevertheless, its applicability has been constrained due to the exclusive availability of MRE mix datasets in Japanese, thereby limiting comprehensive exploration by the global research community. To address this limitation, we introduce a Multilingual MRE mix dataset (MMM) that encompasses 21 sub-datasets in English, Japanese, and Chinese. In this paper, we also propose a method for dataset translation assisted by Large Language Models (LLMs), which significantly reduces the manual annotation time required for dataset construction by leveraging LLMs to translate the original Japanese datasets. Additionally, we have enriched the dataset by incorporating open-domain Named Entity Recognition (NER) and sentence classification tasks. Utilizing this expanded dataset, we developed a unified input-output framework to train an Open-domain Information Extraction Large Language Model (OIELLM). The OIELLM model demonstrates the capability to effectively process novel MMM datasets, exhibiting significant improvements in performance.
@@
MMM: Multilingual Mutual Reinforcement Effect Mix Datasets & Test with Open-domain Information Extraction Large Language Models
@@ 1 Introduction
Figure 1: The Mutual Reinforcement Effect between the labels of Word-level labels and text-level label within a same text.
@@
The Mutual Reinforcement Effect (MRE) Gan et al. (2023b) represents an emerging research direction in multitasking Information Extraction (IE). It primarily explores the interconnections among individual subtasks, thereby enhancing the performance of each task. MRE categorizes IE subtasks into two principal groups: text-level tasks and word-level tasks. For instance, conventional tasks such as sentence classification, text classification, and text sentiment analysis fall under text-level tasks. Conversely, tasks like NER, which involve classifying words, are categorized as word-level tasks.
Figure 1 illustrates a specific example of MRE applied to IE subtasks. The left side of the figure depicts labels for sentence classification, while the right side displays words alongside their corresponding labels extracted from the sentence. Thus, the left side corresponds to text-level tasks, and the right side to word-level tasks. Notably, each sentence features both a text-level label and word-level label entities pair. This differs from typical IE multitasking, which extracts information from various sub-tasks across different texts. Instead, MRE simultaneously performs text-level classification and word-level label-entities pairing on the same text.
We can elucidate the MRE with specific examples. Consider the sentence from Figure 1: ’Giant pandas are mammals, endemic to China.’ This sentence is categorized under ’nature,’ and it includes two annotated label-entities pairs: ’Animal Name: pandas’ and ’Nation: China.’ This exemplifies that when a sentence is classified as relating to nature, it likely contains terms associated with natural elements, such as species names and geographical locations. Similarly, when a sentence is classified as nature, there is a high probability that the sentence as a whole describes nature. The interplay between these two levels of text level classification and word-level entity recognition illustrates the MRE, as each level mutually reinforces understanding and classification accuracy. Expanding upon this, MRE is applicable to various subtasks in IE. For instance, in sentiment analysis, the presence of numerous words with positive sentiment typically indicates that the text overall conveys a positive sentiment. Conversely, if a text is overall negative, it likely contains words with negative sentiment. Integrating tasks at both the text and word levels not only aids in the model’s comprehension of the context but also improves performance across both tasks. This methodology probes whether LLMs can interpret texts similarly to human understanding, by assessing the semantics of individual words and then synthesizing these insights to classify the text as a whole (Gan et al., 2023c).
@@
Figure 2 illustrates the composition of the translated Multilingual Mutual Reinforcement Effect Mix (MMM) Datasets. Notably, the SCPOS subdataset within the original Japanese dataset is substantially larger than others, which is why the figure does not depict the proportional sizes of the subdatasets; instead, it merely labels each subdataset within the MMM dataset. This dataset encompasses three languages and includes seven distinct subdatasets for each language, each corresponding to a specific information extraction task.
Among these, SCNM denotes the Sentence Classification and Named Entity Recognition Mix Dataset. SCPOS, which stands for Sentiment Classification and Part-of-Speech Dataset, involves comprehensive sentiment polarity classification at both the text and word levels. This subdataset is further segmented into four categories: sentiment-related words, adjectives and nouns, solely adjectives, and solely nouns. This categorization facilitates an examination of how words with different lexical properties influence the overall sentiment classification of a text.
Additionally, the TCREE dataset is dedicated to Text Classification and Relation & Event Extraction, leveraging mixed data types for enhanced analytical depth. Lastly, TCONER represents the Open-Domain NER and Text Classification, utilizing an existing open-domain dataset for annotations. This structure allows for a nuanced exploration of text classification and entity recognition across various domains.
This work begins by introducing a novel dataset translation framework. Initially, six MRE mix datasets were translated, and subsequently, the TCONER dataset was expanded. Furthermore, to address the suboptimal performance of existing large language models LLMs on IE tasks, we refined the training methods. We propose a new, streamlined, and efficient input-output scheme that standardizes the handling of all tasks, enhancing the model’s training process. The LLM was then trained using the MMM dataset. The outcomes indicate that the optimized LLM, referred to as OIELLM, surpasses the original training method in performance across various datasets. Moreover, expanding the use of more MRE mix datasets is shown to further enhance the knowledge utilization within the LLM, leading to improved performance across all tasks.
This paper presents several contributions are detailed as follows:
@@
    Development of a Novel Framework for Dataset Translation: We propose a new framework that leverages existing LLMs to translate datasets for underrepresented languages. This framework facilitates the construction of Multilingual MRE Mix Datasets, significantly reducing the reliance on extensive human labor and addressing the issue of dataset scarcity in linguistic research.
@@
    Expansion of MRE Mix Datasets: We have successfully expanded the MRE Mix datasets to include a newly constructed open-domain NER dataset. This addition serves as a critical component in enhancing the comprehensiveness and utility of the MRE Mix datasets, effectively placing the last piece in the puzzle for open-domain applications.
@@
    Optimization of Input and Output Schemes for IE LLMs: Our research optimizes the traditional input and output mechanisms of information extraction LLMs. By employing the specially constructed MMM dataset, we not only reject the standard models trained on generic data but also significantly enhance the performance of LLMs within the domain of information extraction.
@@ 3.2 Construction of TCONER
In the original MRE mix datasets, relation and event extraction tasks are open-domain, implying that the labels are not predefined. However, the label set is limited to only a dozen options. Given this context, we constructed a new dataset, termed TCONER, based on an open-domain Named Entity Recognition (NER) dataset***https://huggingface.co/datasets/Universal-NER/Pile-NER-type?row=0 (Zhou et al., 2023). The labels at the text level in the TCONER dataset are also open-domain. To annotate this dataset, we initially employed the GPT-3.5-Turbo model to assign open-domain text-level labels. Subsequent manual verification and annotation were conducted to ensure accuracy and consistency, resulting in the finalized TCONER dataset. Similarly, we translated the constructed English TCONER dataset using the dataset translation framework. The TCONER dataset was translated into Japanese and Chinese.
@@ 3.3 Results of Datasets Construction
Appendix A Table 2 presents the statistics of the final translation results. Due to the high costs associated with the use of a premium API, we limited our study to 10,000 samples from each of three sub-datasets within SCPOS and the TCONER dataset, which contains 180,000 entries. These 10,000 samples, retained post-translation, proved to be an ample test set. It was observed that there was a greater data loss when translating into Chinese compared to English. This discrepancy may be attributed to the training data predominance of English in OpenAI’s GPT-3.5-Turbo model, resulting in superior performance in English-related tasks. For instance, in the SCNM and TCREE datasets, the Japanese to English translation accuracy exceeded 80%. Conversely, the translation results from English to Chinese in the TCONER dataset were markedly better than those from English to Japanese. This further confirms that GPT-3.5-Turbo exhibits enhanced effectiveness with major languages compared to lesser-used ones.
@@ 5.1 Details of OIELLM Training
We began by selecting baselines: USA-7B (IL + ICL)https://huggingface.co/ganchengguang/USA-7B-instruction-incontext-learning and GIELLM-13B-jphttps://huggingface.co/ganchengguang/GIELLM-13B-jpllm, previously utilized for processing the MRE mixed datasets, served as comparative models. For the foundational architecture of OIELLM, we chose the latest Instruct and Base version of LLaMA3-8B§§§https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. Since LLaMA3 does not offer a 13B version, we incorporated the LLaMA2-13B Touvron et al. (2023) model as well.
@@
OIELLM was fine-tuned using full parameters based on these three models. Training was conducted at BF16 precision, while inference was performed at FP16. The training spanned 3 epochs with a learning rate of 1e-5, utilizing computational resources including three A800 80GB and three RTX 6000 Ada 48GB GPUs, with training durations ranging from 12 to 20 hours. For the training and test sets, Comprehensive statistics on the training and test sets are available in Appendix B.
@@ 5.2 Evaluation
We employed the F1 score as our primary metric for evaluation. Initially, the model’s output was bifurcated into two segments based on the task-specific instruct word: the Text-level Label and the Label-entities pairs. Subsequently, Label-entities pairs were delimited using start-end symbols (i.e., ":", ";"). Each Label-entity pair was treated as an individual element within the set. The F1 score was segmented into three categories: Text-level (TL), Word-level (WL), and ALL. These represent the F1 scores at respective levels and the aggregate F1 score when both levels are accurately predicted in an output. For detailed methodologies, including codes and formulas, please refer to Appendix C.
@@ 6 Results
However, OIELLM’s performance on the TCONER task was suboptimal, which we attribute to insufficient training data. Given that open-domain tasks require extensive and diverse datasets compared to domain-specific tasks, the limited data may have hindered the model’s performance. This area will be a focus of our future research, aiming to understand and improve the data dependencies of OIELLM in open-domain contexts.
@@ 8 Limitations
Appendix A Statistical results of the translated MMM dataset
Dataset SCNM SCPOS: RW SCPOS:
RW Adj & N
Japanese 5343 2000 187528
English 4449 1312 4801
Chinese 3177 1406 3937
Dataset SCPOS: SCPOS: TCREE
Adj N
Japanese 187528 187528 2000
English 9132 5027 1910
Chinese 7413 3920 1491
Language English Japanese Chinese
TCONER 45888 6791 9047
Dataset SCNM SCPOS: RW SCPOS:
RW Adj & N
Japanese 1000 1000 1000
English 1000 500 1000
Chinese 1000 500 1000
Dataset SCPOS: SCPOS: TCREE
Adj N
Japanese 1000 1000 1000
English 1000 1000 500
Chinese 1000 1000 500
Language English Japanese Chinese
TCONER 2000 2000 2000
Table 3: Statistical results of train sets of OIELLM.
Dataset SCNM SCPOS: RW SCPOS:
RW Adj & N
Japanese 4343 1000 186528
English 3449 812 3801
Chinese 2177 906 2937
Dataset SCPOS: SCPOS: TCREE
Adj N
Japanese 186528 186528 1000
English 8132 4027 1410
Chinese 6413 2920 991
Language English Japanese Chinese
TCONER 43888 4791 7047
Table 4: Statistical results of test sets.
Appendix B Statistical Results of Train and Test Dataset in OIELLM
Appendix C Calculate Detail of F1 Score
$F_{1}=2\times\frac{\text{precision}\times\text{recall}}{\text{precision}+\text{recall}}$ (1)
$\text{precision}=\frac{|Real\cap Generated|}{|Generated|}$ (2)
$\text{recall}=\frac{|Real\cap Generated|}{|Real|}$ (3)

### v3

# Point and Line: Multilingual Mutual Reinforcement Effect Mix Information Extraction Datasets

@@ Abstract
The Mutual Reinforcement Effect (MRE) represents a promising avenue in information extraction and multitasking research. Nevertheless, its applicability has been constrained due to the exclusive availability of MRE mix datasets in Japanese, thereby limiting comprehensive exploration by the global research community. To address this limitation, we introduce a Multilingual MRE mix dataset (MMM) that encompasses 21 sub-datasets in English, Japanese, and Chinese. In this paper, we also propose a method for dataset translation assisted by Large Language Models (LLMs), which significantly reduces the manual annotation time required for dataset construction by leveraging LLMs to translate the original Japanese datasets. Additionally, we have enriched the dataset by incorporating open-domain Named Entity Recognition (NER) and sentence classification tasks. Utilizing this expanded dataset, we developed a unified input-output framework to train an Open-domain Information Extraction Large Language Model (OIELLM). The OIELLM model demonstrates the capability to effectively process novel MMM datasets, exhibiting significant improvements in performance. The OIELLM model and datasets is open-source in HuggingFace: GitHub Website***https://ganchengguang.github.io/MRE/
@@
Point and Line: Multilingual Mutual Reinforcement Effect Mix Information Extraction Datasets
@@ 1 Introduction
Figure 1: The Mutual Reinforcement Effect between the labels of Word-level labels and text-level label within a same text. A word-level IE task is a Point, and a text-level IE task is a Line. There is Mutual Reinforcement Effect between the point and the line.
@@
The Mutual Reinforcement Effect (MRE) Gan et al. (2023b) introduces a novel approach in multitasking IE, emphasizing task interconnections to enhance performance. MRE categorizes IE subtasks into text-level tasks (e.g., sentence classification, text sentiment analysis) and word-level tasks (e.g., NER). Unlike conventional IE multitasking, which extracts data from various texts, MRE simultaneously performs text-level classification and word-level label-entities pairing within the same text.
MRE categorizes IE tasks into word-level and text-level tasks, analogous to points and lines. Understanding either part helps reinforce the comprehension of the other. Traditionally, IE subtasks have been studied separately, focusing either on points or lines. MRE, however, is the first approach to integrate these two levels, exploring their interdependencies. This not only enhances the performance of IE subtasks but also has implications for future LLM training. When training data is limited, MRE enables dual-level training of LLMs using a single dataset, maximizing its utility and improving model performance.
Figure 1 illustrates MRE in action. The left side depicts sentence classification labels, while the right side shows words with their corresponding labels, representing text-level and word-level tasks, respectively. For example, the sentence ’Giant pandas are mammals, endemic to China.’ is labeled ’nature’ and contains entity pairs ’Animal Name: pandas’ and ’Nation: China.’ This highlights how text-level classification and word-level entity recognition reinforce each other, improving accuracy.
Similarly, in sentiment analysis, a text with many positive words likely conveys a positive sentiment. Conversely, a negative-text classification indicates the presence of negative words. This interaction mirrors human text comprehension, where meaning is derived from individual words and synthesized into an overall context (Gan et al., 2023c).
@@
Figure 2 shows the composition of the Multilingual Mutual Reinforcement Effect Mix (MMM) Datasets, which include seven subdatasets per language across three languages. Notably, SCPOS, focused on sentiment classification and part-of-speech tagging, is larger than others and thus not depicted proportionally. SCNM involves sentence classification and NER, while TCREE covers text classification, relation, and event extraction. TCONER leverages an open-domain dataset for text classification and NER.
We translated six MRE mix datasets and expanded the TCONER dataset. To improve LLM performance on IE tasks, we refined the training process by introducing a streamlined input-output scheme, standardizing task handling, and training the LLM with the MMM dataset. The resulting optimized model, OIELLM, outperformed previous models on multiple datasets, demonstrating the effectiveness of using expanded MRE mix datasets.
Key contributions include:
@@
    Proposing a framework that uses existing LLMs to translate datasets for underrepresented languages, facilitating multilingual MRE Mix Dataset construction and reducing reliance on manual labor.
@@
    Expanding the MRE Mix datasets with a newly constructed open-domain NER dataset, enhancing coverage and utility for open-domain applications.
@@
    Optimizing IE LLM input-output mechanisms using the MMM dataset, surpassing standard models trained on generic data and improving LLM performance in IE tasks.
@@ 3.2 Construction of TCONER
In the original MRE mix datasets, relation and event extraction tasks are open-domain, implying that the labels are not predefined. However, the label set is limited to only a dozen options. Given this context, we constructed a new dataset, termed TCONER, based on an open-domain Named Entity Recognition (NER) datasethttps://huggingface.co/datasets/Universal-NER/Pile-NER-type?row=0 (Zhou et al., 2023). The labels at the text level in the TCONER dataset are also open-domain. To annotate this dataset, we initially employed the GPT-3.5-Turbo model to assign open-domain text-level labels. Subsequent manual verification and annotation were conducted to ensure accuracy and consistency, resulting in the finalized TCONER dataset. Similarly, we translated the constructed English TCONER dataset using the dataset translation framework. The TCONER dataset was translated into Japanese and Chinese.
@@ 3.3 Results of Datasets Construction
Table 1 presents the statistics of the final translation results. Due to the high costs associated with the use of a premium API, we limited our study to 10,000 samples from each of three sub-datasets within SCPOS and the TCONER dataset, which contains 180,000 entries. These 10,000 samples, retained post-translation, proved to be an ample test set. It was observed that there was a greater data loss when translating into Chinese compared to English. This discrepancy may be attributed to the training data predominance of English in OpenAI’s GPT-3.5-Turbo model, resulting in superior performance in English-related tasks. For instance, in the SCNM and TCREE datasets, the Japanese to English translation accuracy exceeded 80%. Conversely, the translation results from English to Chinese in the TCONER dataset were markedly better than those from English to Japanese. This further confirms that GPT-3.5-Turbo exhibits enhanced effectiveness with major languages compared to lesser-used ones.
Dataset Japanese English Chinese
SCNM 5343 4449 3177
SCPOS: RW 2000 1312 1406
SCPOS: Adj & N 187528 4801 3937
SCPOS: Adj 187528 9132 7413
SCPOS: N 187528 5027 3920
TCREE 2000 1910 1491
TCONER 6791 45888 9047
@@ 5.1 Details of OIELLM Training
We began by selecting baselines: USA-7B (IL + ICL)https://huggingface.co/ganchengguang/USA-7B-instruction-incontext-learning and GIELLM-13B-jp§§§https://huggingface.co/ganchengguang/GIELLM-13B-jpllm, previously utilized for processing the MRE mixed datasets, served as comparative models. For the foundational architecture of OIELLM, we chose the latest Instruct and Base version of LLaMA3-8Bhttps://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. Since LLaMA3 does not offer a 13B version, we incorporated the LLaMA2-13B Touvron et al. (2023) model as well.
@@
OIELLM was fine-tuned using full parameters based on these three models. Training was conducted at BF16 precision, while inference was performed at FP16. The training spanned 3 epochs with a learning rate of 1e-5, utilizing computational resources including three A800 80GB and three RTX 6000 Ada 48GB GPUs, with training durations ranging from 12 to 20 hours. For the training and test sets, Comprehensive statistics on the training and test sets are available in Table 3, 4.
Dataset Japanese English Chinese
SCNM 1000 1000 1000
SCPOS: RW 1000 500 500
SCPOS: Adj & N 1000 1000 1000
SCPOS: Adj 1000 1000 1000
SCPOS: N 1000 1000 1000
TCREE 1000 500 500
TCONER 2000 2000 2000
Table 3: Statistical results of train sets of OIELLM.
Dataset Japanese English Chinese
SCNM 4343 3449 2177
SCPOS: RW 1000 812 906
SCPOS: Adj & N 186528 3801 2937
SCPOS: Adj 186528 8132 6413
SCPOS: N 186528 4027 2920
TCREE 1000 1410 991
TCONER 4791 43888 7047
Table 4: Statistical results of test sets.
6 Statistical Results of Train and Test Dataset in OIELLM
@@ 6.1 Evaluation
We employed the F1 score as our primary metric for evaluation. Initially, the model’s output was bifurcated into two segments based on the task-specific instruct word: the Text-level Label and the Label-entities pairs. Subsequently, Label-entities pairs were delimited using start-end symbols (i.e., ":", ";"). Each Label-entity pair was treated as an individual element within the set. The F1 score was segmented into three categories: Text-level (TL), Word-level (WL), and ALL. These represent the F1 scores at respective levels and the aggregate F1 score when both levels are accurately predicted in an output. The formulas for calculating the F1 score from the set of segmented label-entities pairs are shown in Equations (1) - (3) below. See Appendix A for specific pseudo-code for the entire evaluation.
$\text{Precision}=\frac{|Real\cap Generated|}{|Generated|}$ (1)
$\text{Recall}=\frac{|Real\cap Generated|}{|Real|}$ (2)
$F_{1}\text{ Score}=2\times\frac{\text{precision}\times\text{recall}}{\text{precision}+\text{recall}}$ (3)
@@ 7 Results
However, OIELLM’s performance on the TCONER task was suboptimal, which we attribute to insufficient training data. Given that open-domain tasks require extensive and diverse datasets compared to domain-specific tasks, the limited data may have hindered the model’s performance. This area will be a focus of our future research, aiming to understand and improve the data dependencies of OIELLM in open-domain contexts. Due to the high cost of accessing GPT-4/o, we conducted experiments on MMM datasets using GPT-3.5-Turbo only. Despite tailoring prompt templates to align with each dataset’s specific features, GPT-3.5-Turbo demonstrated subpar performance. This can be attributed to the model’s limited suitability for tasks requiring the simultaneous extraction of multiple label-entity pairs. It frequently generates redundant fields and extraneous characters, resulting in a lower F1 score. These limitations underscore the necessity of developing an OIELLM specifically trained for this task.
@@ 9 Limitations
Appendix A Evaluation Pseud Code
@@
Appendix B Case Study of Input and Output Format with OIELLM in MRE mix datasets
Figure 6: The input and output format example with OIELLM in Japanese MRE mix datasets.
Figure 7: The input and output format example with OIELLM in English MRE mix datasets.
Figure 8: The input and output format example with OIELLM in Chinese MRE mix datasets.